### v2

# Day-Ahead Electricity Price Forecasting Using Merit-Order Curves Time Series

Abstract
We introduce a general, simple, and computationally efficient framework for predicting day-ahead supply and demand merit-order curves, from which both point and probabilistic electricity price forecasts can be derived. We conduct a rigorous empirical comparison of price forecasting performance between the proposed curve-based model, i.e., derived from predicted merit-order curves, and state-of-the-art price-based models that directly forecast the clearing price, using data from the Italian day-ahead market over the 2023-2024 period. Our results show that the proposed curve-based approach significantly improves both point and probabilistic price forecasting accuracy relative to price-based approaches, with average gains of approximately 5%, and improvements of up to 10% during mid-day hours, when prices occasionally drop due to high renewable generation and low demand.
keywords:
energy forecasting , electricity , price forecasting , day-ahead market , merit-order curves , functional data analysis , multivariate time series , vector autoregression models , probability forecasting[mox]organization=MOX, Department of Mathematics, Politecnico di Milano,addressline=Piazza Leonardo da Vinci 32, city=Milano, postcode=20133, state=MI, country=Italy
[deng]organization=Department of Energy, Politecnico di Milano,addressline=Via Lambruschini 4a, city=Milano, postcode=20156, state=MI, country=Italy
1 Introduction
1.1 Context
The rising integration of non-programmable renewable energy sources and the growing uncertainty surrounding fossil fuel supply have made electricity markets more volatile and difficult to forecast. As a result, the development of sophisticated and reliable market forecasting tools has become even more essential for all actors of the electric power industry.
Most wholesale electricity trading in Europe takes place on day-ahead markets, where producers and retailers exchange power one day before delivery. These markets generally operate under a double uniform-price (pay-as-clear) auction mechanism, where the intersection of the supply and demand merit-order curves sets the market clearing price for all accepted offers and bids. These curves are formed from the aggregation of individual supply offers and demand bids following the merit order, i.e. increasing price for supply and decreasing price for demand (Fig. 1). Offers (or bids) priced below (resp. above) the clearing price are accepted111The intersection point may split a supply offer or demand bid which in that case would be partially accepted while the remainder are rejected. For market participants, accurately estimating the clearing price is therefore sufficient to anticipate whether an offer/bid will be accepted and at which price.
Because of this single-price property, a large body of research has focused on forecasting the day-ahead clearing price itself, a discipline known as electricity price forecasting (EPF). The consensus is that data-driven approaches based on historical prices tend to outperform fundamental (or structural) ones which try to model exact market mechanisms, especially for short-term hourly forecasts (Weron, 2014). In particular, a highly competitive day-ahead EPF benchmark based on regularized autoregressive models and feedforward neural networks was consolidated by Lago et al. (2021). This framework is still state-of-the-art and has been widely employed in recent studies (Olivares et al., 2023; Lipiecki et al., 2024; O’Connor et al., 2025; Cerasa and Zani, 2025).
Nevertheless, the problem of forecasting the entire merit-order curves underlying price formation is gaining momentum in the academic community (Petropoulos et al., 2022). Main contributions on this topic can be found in Table 1. The seminal work of Ziel and Steinert (2016) demonstrated the capacity of ”data-driven fundamental” approaches based on historical auction data and their potential for generating competitive clearing price forecasts. Forecasting entire curves offers several advantages. For participants with market power, it enables precise quantification of how marginal changes in offers may influence the clearing price and allows realistic market simulations from which optimal bidding strategies can be derived (Pelagatti, 2013). For price-takers and other actors exposed to the day-ahead price, curve-based models offer a more comprehensive view of market results, enhances the interpretability of price predictions, and may improve the quantification of price predictions’ uncertainty222This latter aspect is investigated in this paper.. Additionally, complex nonlinear dynamics observed at the price level may become more tractable when modeled at the curve level.
Work Model typology Dim. reduction EPF Probabilistic EPF Pelagatti (2013) VAR FPCA Ziel and Steinert (2016) VAR Discretization ✓ ✓ Canale and Vantini (2016) Functional AR ✗ Mestre et al. (2020) Functional AR ✗ Kulakov (2020) VAR Discretization ✓ Shah and Lisi (2020) Functional AR ✗ ✓ ✓ Guo et al. (2021) Neural Network PCA Soloviova and Vargiolu (2021) Functional AR ✗ ✓ Yıldırım et al. (2023) State-space Parametrization Ciarreta et al. (2023) VAR Parametrization ✓ Tang et al. (2024) Neural Network PCA, NMF, SDL Vivó and Alonso (2024) Neural Network ✗ Ghelasi and Ziel (2024) VAR Discretization Li et al. (2024) k-NN ✗ Ghelasi and Ziel (2025) Fundamental ✗ ✓ Sinha and Lucheroni (2025) Neural Network Autoencoder Li et al. (2025) ARIMA, Functional AR FPCA ✓
1.2 Knowledge gap
Given the absence of a unified evaluation framework – as in EPF – and the scarcity of direct and fair comparisons between proposed approaches, identifying top-performing merit-order curves forecasting techniques remains challenging. In particular, aside from a few direct extensions (Kulakov, 2020; Ghelasi and Ziel, 2024) and despite a large number of subsequent contributions, no alternative method has been conclusively shown to outperform the original approach of Ziel and Steinert (2016). We identified only one study, Yıldırım et al. (2023), that provides a direct comparison, though limited to supply curve forecast accuracy. Their results suggest that their method and that of Ziel and Steinert perform similarly. A comparison of curve forecast performance is also reported in Sinha and Lucheroni (2025), although the authors state that they employ a modified version of Ziel and Steinert’s model without providing further details333Additionally, we are somewhat skeptical of their results as all considered benchmarks – including Ziel and Steinert’s model – are reported to perform far worse than a naive method.. As a result, Ziel and Steinert’s framework remains the de facto state-of-the-art, and whether it can be consistently improved remains an open question.
Beyond the problem of curve forecasting itself, an essential question remains unanswered: are merit-order curve forecasting models competitive with state-of-the-art price-centric EPF models in terms of clearing price prediction accuracy?
While Ziel and Steinert demonstrated ten years ago that their approach outperformed existing price-based models for point forecasting, this result may no longer hold in light of recent advances in EPF. Since then, several studies have assessed the effectiveness of curve-based approaches for clearing price prediction by comparing them with price-based methods; however, the latter are often overly simplistic and not representative of top-performing EPF models. In addition, the rigorous forecast evaluation practices recommended by Lago et al. (2021), such as daily recalibration, are seldom followed. Furthermore, none of the post-2021 contributions compare curve-based price forecasts against the benchmark models of Lago et al., despite their availability through an open-access software implementation (the epftoolbox Python package). For instance, Shah and Lisi (2020) compare their curve-based forecasts only against basic autoregressive models without exogenous covariates, even though variables such as load and renewable generation forecasts are known to be essential for achieving top performance (Weron, 2014). Similar limitations apply to Soloviova and Vargiolu (2021), Ciarreta et al. (2023), and Li et al. (2025). Ghelasi and Ziel (2025) consider an “expert” autoregressive model with exogenous covariates, but do not include regularized, parameter-rich autoregressive models with cross-hour dependence, despite evidence that such models consistently outperform expert ARX specifications (Ziel and Weron, 2018).
1.3 Contributions
In this context, this work has two main objectives:
1.
Introduce a general, simple, and computationally efficient framework for predicting day-ahead supply and demand merit-order curves, from which both point and probabilistic clearing price forecasts can be derived.
2.
Rigorously compare the clearing price forecasting performance of the proposed curve-based models, i.e. derived from merit-order curves predictions, with that of state-of-the-art price-based models, which directly model the clearing price.
Our framework builds on Ziel and Steinert (2016) but eliminates the need for curves discretization by preserving their functional form. Specifically, we leverage functional principal component analysis to efficiently represent a pair of supply and demand curves in a vector space and employ regularized vector autoregressive models inspired by EPF literature for their prediction. We compare this functional approach with the original method of Ziel and Steinert and top-performing price-based point and probabilistic EPF models, on the Italian day-ahead market (GME), during the 2023-2024 period.
2 Methodology
2.1 Curves definition
Formally, a merit-order curve represents price as a function of quantity and is defined from a set of supply offers (demand bids, respectively) characterized by a price and a quantity sorted following the merit-order, i.e., increasing order of price for the supply curve, decreasing order of price for the demand curve. Mathematically, a merit-order curve can be defined as follows:
where we set , for notational convenience. For supply (or demand), is said to be the marginal price associated to a total offered (resp. demanded) quantity .
From the definition we can notice that merit-order curves are non-continuous monotone (increasing for supply and decreasing for demand) step functions. A direct consequence is that admits an inverse which is itself a monotone step function. It is interesting to note that this inverse has an easier mathematical notation and interpretation. Indeed, its expression matches (for supply curves) the unnormalized444Given the ’s don’t sum to 1. cumulative distribution function (cdf) of a discrete distribution over the support . For demand, the only difference is that it is the reverse cdf (or survival function). Using the superscript for supply and for demand, we have:
That is, is the quantity that is offered (or demanded) at a price below (resp. above) . Note that the definition above does not require the sequence of offers (or bids) to be price-ordered. It is more intuitive to work with the quantity function as it represents a more familiar mathematical object than the price function . The quantity function simply tells us about the total offered quantity and how this offered quantity is distributed on the price domain. It also conveniently changes the perspective on the curves variability which is mainly attributed to their inelastic component: seasonal power demand variations for demand curves and variable renewable energy generation for supply curves.
From now on, unless otherwise specified, references to supply or demand curves will refer to the quantity function.
2.2 Curves representation
The goal is to model and forecast the time series of curves and over a time period composed of hourly intervals. Functional data analysis (Ramsay and Silverman, 2005; Ferraty and Vieu, 2006) is a framework that extends traditional statistical methods to handle data that vary over a continuum, such as curves, surfaces, or other functional forms. In order to forecast these functional time series, we follow the guidance of Aue et al. (2015) that recommend the use of functional principal component analysis (FPCA) to derive an uncorrelated vector representation of the curves and apply standard multivariate time series forecasting methods to this representation, finally obtaining a curves forecast using the Karhunen-Loève (KL) expansion (Horváth and Kokoszka, 2012, pp. 37–43). FPCA is a dimensionality reduction technique for functional data, extending classical PCA to infinite-dimensional spaces. It decomposes a set of observed functions into an orthonormal basis that captures the dominant modes of variations in the data. This decomposition yields the truncated KL expansion:
where is the mean function, are the first functional principal components (FPCs) and are the scores of observation on the first FPCs. Intuitively, the FPCs can be seen as specific features common to all curves while the scores are measures of how pronounced are these features in a specific curve. The perfect equality holds when but in practice a satisfying approximation can be found with a finite , typically chosen – as in traditional PCA – as the elbow point of the scree plot or such that the ratio of explained variance surpasses a certain threshold (Johnson and Wichern, 2007, pp. 444–447).
Since FDA assumes smooth functions, the estimated FPCs should also be smooth functions. This smoothness can be achieved either by incorporating a roughness penalty directly into the FPCA estimation procedure or by pre-smoothing the functional data curves prior to applying FPCA. We adopt the latter approach for computational simplicity, utilizing kernel smoothing with the Nadaraya-Watson kernel estimator (Wasserman, 2006, p. 71). The bandwidth parameter of the kernel function, which controls the level of smoothness, is chosen using the generalized cross-validation criterion (Ramsay and Silverman, 2005, p. 97).
In our case, we would like to account for possible dependence between the supply and demand curves time series. To do so, we perform FPCA separately for the supply and demand curves and by concatenating the scores for the supply curves with the scores of the demand curve, we get a -dimensional vector representation of the paired functional series , with
We forecast with some multivariate forecasting model (detailed in the next section), with possibly a set of exogenous covariates . We then inverse-transform any -step-ahead forecast using the truncated KL expansion to get a forecast for the curves pair .
It can happen that curves predictions are not perfectly monotonic. Though preliminary experiments suggested the occurence of little deviations only and low impact on clearing price predictions, it is still desirable to respect this structural constraint. To do so, we post-process the curves predictions using isotonic regression with the pool-adjacent violators algorithm (Leeuw et al., 2009). Being a highly efficient algorithm, it has a negligible impact on computation times.
2.3 Forecasting the curves’ vector representation
We consider here the day-ahead forecasting problem, regarding the vector hourly time series as 24 separate daily time series, one for each hour of the day. This is common practice in EPF as the 24 hours of the next day are simultaneously settled the day before. In addition, market dynamics vary a lot depending on the hour of the day, justifying to treat them separately. For a detailed motivation of this choice, the reader can refer to Ziel and Weron (2018). We hence change the time indexing of to consider the value at day and hour , and we solve 24 one-step-ahead forecasting problems.
We consider four variants of the popular parameter-rich ARX model estimated with LASSO widely used in the context of electricity price forecasting (Ziel, 2016; Uniejewski et al., 2016; Ziel and Weron, 2018; Lago et al., 2021; Uniejewski, 2024). First this framework was shown to be highly performing for price forecasting and second, it was used in Ziel and Steinert (2016). Specifically, we consider two univariate models treating each component of separately as suggested by Hyndman and Shang (2009) – the concurrent ARX and the full ARX (equations 1 and 2) – and two multivariate models jointly modeling the vector , the concurrent VARX and the full VARX (equations 3 and 4). The term ”concurrent” means the target variable at hour is influenced only by past values from the same hour . Conversely, ”full” means the variable is also influenced by past values observed at different hours. The comparison between the univariate and the multivariate approaches allow to test the added value of modeling the cross-dependence between scores while the comparison between the concurrent and full approaches allow to test the added value of modeling the cross-dependence between hours.
The *AR* part refers to the fact that lagged values are used to predict future values. In our case, similarly to state-of-the-art EPF models (Lago et al., 2021) we include lags 1, 2, 3 and 7, considering therefore information up to 7 days before. Additionally, each model has the suffix *X which means that it makes use of exogenous variables which refer to day and available on (i.e., they are known in anticipation or they are themselves one-day-ahead forecasts). This set of exogenous variables is represented by the -dimensional vector . Again, like Lago et al. (2021), we also include lags 1 and 7 of these exogenous variables, and a three-dimensional vector of dummy variables flagging the day type: Mondays, Working days (from Tuesday to Friday), Saturdays and Holidays (Sundays and bank holidays)555Note this is slightly different from Lago et al. (2021) who consider one dummy for each day of the week but similar to Mestre et al. (2020) and addresses the calendar effects issues raised by Ziel and Steinert (2016)..
Finally, still following Lago et al.’s guidance, the -regularization parameter is selected as that minimizing the Akaike information criterion (AIC), exploiting the least angle regression (LARS) algorithm (Efron et al., 2004) for the parameter search. Once is selected, the model is estimated using the traditional coordinate descent algorithm (Tibshirani, 1996).
Concurrent ARX (ARX)
The concurrent ARX models each component of as:
where are the autoregressive coefficients and and are the -dimensional and 3-dimensional vectors of (lagged) exogenous and dummy variables coefficients, respectively.
Full ARX (fARX)
The full ARX , very similar to its homonym in Uniejewski et al. (2016), models each component of as:
where is the linear effect of the lagged hour on the hour .
Concurrent VARX (VARX)
The concurrent VARX models the full vector as:
Where are matrices of autoregressive coefficients and and are the and matrices of (lagged) exogenous and dummy variables coefficients, respectively. Contrarily to the two previous models, we allow for cross-dependence – described by the off-diagonal coefficients of – between the components of .
Full VARX (fVARX)
The full VARX models the full vector as:
Where we impose the restriction that the off-diagonal terms of must be zero whenever , that is, cross-dependence is allowed between hours, but only within the same component666In that sense, ”semi-full VARX” could be a more precise denomination but we keep ”full VARX” for simplicity.. This restriction is imposed to have a manageable number of parameters. Indeed, without this condition we would have parameters for each component of while with this condition the number of parameters is only . For instance, for and , the condition implies 147 parameters instead of 975.
2.4 Clearing price point and probabilistic forecasting
The clearing price point prediction is simply obtained by determining the zero777Computationally speaking, we take the linear interpolation of the two prices concerned by the sign change. of the function .
Regarding the probabilistic prediction, the goal is to forecast the quantiles of the conditional distribution of the price at day and hour given all information available at . The strategy we adopt is rather straightforward and based on Monte Carlo simulations:
1.
Estimate the predictive distribution of nonparametrically using a residual-based bootstrap on a rolling calibration window
2.
Simulate a vector from that distribution
3.
Inverse transform to obtain
4.
Find the clearing price
5.
Repeat (2), (3) and (4) times
6.
Take the quantiles of the empirical clearing price distribution obtained.
Details for step (1) can be found in A. The choice of the calibration window is a classic trade-off as short windows have the advantage of capturing potential non-stationarity and correcting eventual forecasts bias while long windows will allow for a more accurate estimation of the errors distribution. To get ”the best of both worlds”, we proceed as in Lipiecki et al. (2024): We consider four different window sizes: 28 days (4 weeks), 56 days (8 weeks), 91 days (13 weeks) and 182 days (26 weeks) yielding four cdf estimates . We then compute an ensemble prediction considering vertical (probability) averaging (Lichtendahl et al., 2013):
2.5 Benchmark curves and price forecasting models
Naive model
The naive model (Naive) works as follows:
Ziel and Steinert (2016)
While our approach differs from that of Ziel and Steinert (2016) regarding the curves vector representation, the multivariate forecasting techniques applied to that representation are essentially the same. For that reason, to guarantee a fair comparison between the two approaches, we apply the models detailed in the previous section to (i) our vector of supply and demand FPCA scores (ii) the vector of supply and demand price class volumes obtained with the method described in Ziel and Steinert (2016). We invite the reader to refer to the authors’ paper for a rigorous and detailed description of this procedure (as well as for curves reconstruction), which we refer to as Ziel-Steinert Transformation (ZST), as opposed to our functional principal component analysis (FPCA).
Briefly, ZST consists of taking the first-order differences of discrete evaluations of the quantity curve on a specific price grid. This price grid is constructed by defining an equispaced quantity (or volume) grid and transforming it into a non-uniform price grid via the mean price curve . For inverse transformation, the differences are cumulatively summed to recover the discrete values and the continuous curves are reconstructed using interpolation based on the mean quantity curve. Fig. 2 shows an example of a transformed and reconstructed supply-demand curves pair.
Price-based point forecasting models
We consider three benchmark price-based clearing price forecasting models: ARX, fARX and the Lasso-Estimated AutoRegressive model (LEAR) of Lago et al. (2021). The first two are already described in the previous section (equations 1 and 2). When both curve-based and price-based models are involved, to distinguish the clearing price forecasting model deriving from the curves model (curve-based) from that directly forecasting the price (price-based), we prefix the curve-based models with the curves representation used, e.g, FPCA-ARX or ZST-ARX.
Finally the LEAR model (equation 6) is very similar to fARX with the difference that it adds cross-hour dependence in the exogenous variables effects and accepts only two of them (usually the load and combined wind-solar generation day-ahead forecasts):
As in Lago et al. (2021), for all three models, prices are transformed with median-based scaling and the area hyperbolic sine variance stabilizing transformation.
Price-based probabilistic forecasting models
We test the three (price-based) clearing price probabilistic forecasting methods of Lipiecki et al. (2024), using the Julia package PostForecasts.jl (Lipiecki and Weron, 2025) developed by the authors: Quantile Regression Machine (QRM), Conformal Prediction (CP), Isotonic Distributional Regression (IDR). We also include the naive Normal (N) benchmark888Note that differently from Lipiecki et al. (2024) paper, we consider the version which does not suppose zero-mean errors and therefore accounts for potential forecasts bias.. As these methods are based on post-processing of point predictions, we use the best performing price-based model as a point predictor. For more information, we invite the reader to directly refer to Lipiecki et al. (2024).
2.6 Forecasts evaluation
As suggested in Ramsay and Silverman (2005), we measure curves forecasting accuracy with the squared correlation function:
which represents the proportion of variance explained by the model across the price domain, and the average squared correlation:
Additionally, to test whether a model significantly outperforms another, we run the Diebold-Mariano (DM) test (Diebold and Mariano, 1995) on the -norm of daily functional forecast errors considering (i) the average error across the 24 hours (ii) separate tests run on each of the 24 hours999Considering daily forecast errors avoids the problem of the strong intraday error autocorrelation, typical in day-ahead price forecasting, which would violate DM tests assumptions (Weron, 2014).. Two one-sided tests are run for each pair of models.
We evaluate clearing price point predictions with the mean absolute error (MAE), the root mean squared error (RMSE) and the relative MAE (rMAE) as defined in Lago et al. (2021).
Price probabilistic predictions are evaluated using the continuous-ranked probability score (CRPS):
where is the realized price at time and is the estimated conditional cdf we want to evaluate.
Similarly to the curves forecasts, we perform DM tests for both point and probabilistic forecasts considering average and per-hour daily errors. For point forecasting we consider the absolute () errors as in Lago et al. (2021) while we use the CRPS, as was done in Nowotarski and Weron (2018), for probabilistic forecasting.
2.7 Daily recalibration
As advised by Lago et al. (2021) and in order to simulate a live forecasting setting, the entire modeling pipeline is recalibrated on a daily basis using a rolling window of 364 days (52 weeks), since year appears to the best single-window choice for EPF according to Marcjasz et al. (2018) and Hubicka et al. (2019). This means that for every day of the testing period:
1.
We extract the vector representation of the curves pairs observed on
2.
We fit the multivariate forecasting model on these vector observations
3.
We predict the 24 curves vector representations for day
4.
We inverse transform the predicted vectors to obtain the 24 curves pairs forecasts
Regarding step (1) this implies that the FPCA or ZST transformation is re-run from scratch every day. This dynamic ”representation learning” procedure is clearly preferable to a time-invariant representation learned on a fixed initial time window but adds a layer of complexity: We must choose the supply and demand dimensions in an automatic manner.
For FPCA, we combine the elbow and threshold methods (see section 2.2) by taking the maximum of the two suggested number of FPCs to retain, considering a ratio of 99% for the threshold method. While automating the latter is straightforward, the elbow method relies on a qualitative visual assessment of the scree plot and requires therefore a more sophisticated approach. To automatically find this elbow point in the recalibration procedure, we use the popular knee-point detection method of Satopaa et al. (2011).
For ZST, Ziel and Steinert (2016) do not provide discussion on their choice for the number of price classes to consider. Since our goal is to compare the efficiency of the two curves representation strategies, we should ideally match it on a daily basis with the dimension picked by FPCA. For simplicity though, we take fixed and equal to their maximum value observed across the FPCA recalibration procedure, leaving therefore a slight advantage to ZST.
Note that this same daily recalibration procedure applies to the price-based models.
3 Results & Discussion
We analyze the day-ahead electricity market (MGP) of the Italian Power Exchange (IPEX) over the 2023-2024 period. The testing period covers the whole year 2024 (366 days). Note that probabilistic price predictions require at least 182 days of out-of-sample point predictions so the testing period for probabilistic forecasting starts on July 1st, 2024, covering a period of 6 months. The quantity curves are built according to the procedure described in section 2.1, using the DomandaOfferta dataset retrieved from the FTP server of Gestore del Mercato Elettrico (GME) – the operator of IPEX – which contains all supply offers and demand bids submitted on the MGP. Additionally, we consider a total of 4 exogenous variables for both curves and price forecasts which are similar to those of Mestre et al. (2020) and commonly chosen in EPF: load forecast, wind and solar generation forecast, and forecasts of transfer capacity with neighboring zones (France and Switzerland). Additional details on curves and exogenous data are provided in B.
The analysis of FPCA results and an empirical assessment of computational efficiency are left in C and D, respectively, while an example of curves and clearing price forecast for each of the nine curve-based models is pictured in E (Fig. 13).
3.1 Curves forecasting
[…]
The average squared correlation and the squared correlation function of the nine curves forecasting models are presented in Table 2 and Fig. 3, respectively. The naive model is clearly outperformed by all the other models for both sides. For demand, all models (except naive) perform almost equally: This is very likely due to the low demand elasticity on the Italian day-ahead market meaning that the quantity demand curve is more or less constant equal to the load forecast – predictor to which all models have access to. Concerning supply, the FPCA representation (FPCA) clearly outperforms that of Ziel and Steinert (2016) (ZST), while the selected variant of the multivariate forecasting model does not show much impact, especially for FPCA. Looking at the entire squared correlation curve, we notice the performance drop of ZST models in the high prices range, which partly explains their lower average .
3.2 Price point forecasting
| MAE | rMAE | RMSE | ||
[…]
The global clearing price forecasting performance is reported on Table 3 with the associated DM tests in Fig. 4(a), for the four price-based and eight curve-based models. All models perform well compared to the naive approach. The DM tests reveal that all ZST models are significantly outperformed by all FPCA, regardless of the multivariate model variant considered. Besides, none of ZST models significantly outperform any of the price-based models, while all of them but one are significantly outperformed by at least one of the price-based models. The FPCA curve-based models instead are the best performers as all of them significantly beat all ZST and price-based approaches101010At the 5% level except for the univariate FPCA models vs fARX for which it is 10%.. A key result is that the consideration of the cross-scores dependence (VARX models) appear to have a significant (positive) impact while the cross-hour dependence (full models) does not show any improvement, whether it be for ZST or FPCA. Instead, for the price-based models, the cross-hour dependence is necessary to reach top performance.
Note the large performance gap between ZST-ARX and ZST-VARX (almost 1 €/MWh) compared to that observed between FPCA-ARX and FPCA-VARX (around 0.20€/MWh). This observation confirms a priori beliefs on the efficiency of the functional vs discretization representation method for capturing the statistical dependence that may exist between different portions of curves domain.
Fig. 5 shows the hour-level performance of each category’s best model: fARX for price-based, ZST-VARX for ZST curve-based and FPCA-VARX for FPCA curve-based, while DM tests performed at the hour-level can be found in Fig. 4(b). We understand that the FPCA model performs particularly well around mid-day, between 12 and 16, matching the period of simultaneous low-demand and high RES generation occasioning price drops, while the price-based model shows a slight advantage – confirmed by DM tests – during the night.
To understand more precisely the difference between the price-based and curve-based forecasts and visualise performance across the price range, we can refer to Fig. 6 where predicted prices are plotted against true prices for each model. Without much surprise, largest errors occur at the tails of the distribution, particularly on the left-side i.e. for prices below 100€/MWh. Note the two opposed biases between price-based and FPCA curve-based models as the first systematically overestimate prices below €/MWh while the second tend to underestimate them. This observation suggests that considering an ensemble of price-based and curve-based models is likely to provide significant improvements.
3.3 Price probabilistic forecasting
| 28D | 56D | 91D | 182D | Avg. | ||
[…]
The average CRPS of the clearing price probabilistic forecasting models are reported in Table 4 with the associated DM tests in Fig. 8. Again, all models outperform the naive strategy but the curve-based models based on ZST representation do not perform well compared to that based on FPCA and price-based models. While FPCA curve-based models perform globally better than price-based approaches, fARX-QRM strongly benefits from the ensemble averaging effect and only two of the curve-based models significantly outperform it at the 10% level regarding daily average errors. The top performers for point forecasting were the multivariate FPCA models, but for probabilistic forecasting, results suggest a slight advantage for univariate versions.
Fig. 7 shows the hour-level average CRPS of the best models of each category. As for point forecasting, the best curve-based model performs better in the early afternoon between 12 and 16 than the price-based approach. The performances are nearly equal for the rest of the day. Reliability and sharpness assessment can be found in F.
4 Conclusion
This work introduced a general framework for forecasting day-ahead merit-order curves, leveraging functional principal component analysis to efficiently represent a pair of supply and demand curves in a vector space and employing regularized vector autoregressive models for their prediction. The application to the Italian day-ahead market during the 2023-2024 period not only demonstrated the method’s effectiveness in forecasting merit-order curves but also in producing highly accurate clearing price point and probabilistic forecasts.
We tested four variations of our model, each treating the hourly time series as 24 independent daily time series – one for each hour – and differing in whether they accounted for (i) cross-dependence between hours and (ii) cross-dependence between components of the vector representation. Our findings indicate that accounting for cross-component dependence slightly improves clearing price forecasting performance, while cross-hour dependence does not yield significant gains.
Our curve-based model was rigorously compared with that of Ziel and Steinert (2016) – which share the same multivariate forecasting framework and differs only in the curves representation strategy – and state-of-the-art price-based models for clearing price forecasting. The results show that considering a functional principal component representation brings consequent improvements over a discrete representation and that our curve-based model – even in its simplest form – significantly improves point and probabilistic clearing price forecasting performance, especially around mid-day when price occasionally drops due to high renewable generation and low demand. Specifically, our curve-based model reduces the mean absolute error by 5% and the continuous ranked probability score by 4%. If we restrict the comparison to the 10:00-17:00 hour range, the reduction is of 10% and 7%, respectively.
Additional analyses could examine the effect of alternative curves representations, the vector space dimension, the use of a single versus hour-specific representation and the calibration window size, on both curves and clearing price forecasting performance. Besides, more complex non-linear models such as feedforward neural networks could be considered to jointly predict the 24 vector representations within a single model. Finally, the recent switch to 15-min intervals in European day-ahead markets, making no longer 24 but 96 curves pairs to predict, could justify a functional rather than multivariate perspective to capture intraday dependence. These aspects will be explored in future research.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the authors used OpenAI ChatGPT and Google Gemini in order to obtain suggestions for improving writing fluency and clarity. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article.
Appendix A Multivariate predictive distribution estimation
We estimate the multivariate distribution of the -dimensional forecast error vector . To robustly estimate this distribution which we assume stationary within the calibration window, we consider hour-dependent variances for the marginals but a shared correlation structure: We assume the errors for a given hour are independent and identically distributed (i.i.d.) according to some -dimensional distribution with mean and a covariance matrix which can be decomposed as follows:
where is the diagonal matrix containing the hour-specific marginals variances and the correlation matrix, assumed identical for all hours. Hence, by bootstrapping the marginal-standardized residuals
across all hours, we obtain simulations which can be transformed in simulations of :
This modeling choice allows us to account for possible hour-dependent uncertainty without having to estimate 24 distinct distributions.
Appendix B Details on curves and exogenous data
Relaxation of transmission constraints
Around 35% of the time, the country-level economic optimum is not feasible due to transmission capacity constraints, which means that the market cannot be cleared from the country-level supply and demand curves. In this case, GME runs an optimization procedure, known as locational marginal pricing (LMP), which splits the entire country-level market pool in a minimum number of ”subpools” of market zones such that, in each subpool, the economic optimum is feasible. As a result, there can be several pairs of supply and demand curves – hence several market clearing prices – for a single hour, one for each subpool. To keep the analysis simple, we did as if the country-level economic optimum was always feasible and built the country-level supply and demand curves for any hour, ignoring the problem of transmission congestion. As a result, the constructed curves reflect the true market clearing 65% of the time, the remaining 35% corresponding to a “virtual” market clearing only. Note the same curves construction procedure was considered in Mestre et al. (2020). The associated market clearing price is what GME calls the national price without constraints, and though not always being the exact market price for the different Italian bidding zones, is still a key monitored market indicator.
Market coupling
Italy is integrated into the European electricity market through market coupling with several neighboring countries which are France, Austria, Slovenia and Greece. Market coupling allocates cross-border trading capacity in order to harmonize European prices. It therefore imposes the cross-border exchanges between coupled countries. As a result, the purchased quantity in MGP does not necessarily equal the sold quantity: if Italy imports more than it exports, the (domestic) sold quantity will be lower than the purchased quantity because a part of the demand is already satisfied by the imports. Conversely, if Italy exports more than it imports, the sold quantity will be higher than the purchased quantity because part of the supply must satisfy the exports. As a consequence, the market clearing price is not rigorously at the intersection of the supply and demand curves. However, everything happens as if imports were entering in the supply offers at the minimum price – such that the underlying offer is necessarily accepted – causing a right shift of the price supply curve, while foreign exports enter in the demand bids at the maximum price – such that the underlying bid is necessarily accepted – causing a right shift of the price demand curve. Therefore by incorporating the imports and exports in the supply and demand curves, the market clearing is found exactly at the intersection. In our case, the import and export quantities within the scope of market coupling are retrieved from the MarketCoupling GME dataset and the difference between imports and exports is added to the quantity supply curve, which is equivalent to adding them separately to both curves (from the point of view of market clearing).
Restriction of the curves domain
A last observation regards the domain on which the quantity curves are analyzed. The entire price domain ranges from -500€ to 3000€. Obviously, as stressed in the introduction, the part of that domain where the intersection point has negligible chances to occur is of lower interest, may be subject to arbitrary variability that has no impact of the market outcomes and could just add undesired noise in the model. Therefore, we decided to restrict the quantity curves to a price domain that contains at least all prices in the 2023-2024 period with a generous margin: indeed, we restrict the curves to the domain comprised between 0€/MWh and 300€/MWh.
Three example pairs of supply and demand curves of the dataset are showed on Fig. 9.
Exogenous predictors
We consider a total of 4 exogenous variables for both curves and price forecasts which are similar to those of Mestre et al. (2020) and commonly chosen in EPF:
1.
Day-ahead load forecast (1 variable): Country-level day-ahead load forecast obtained from the ENTSO-E Transparency Platform (ENTSO-E, 2025).
2.
Day-ahead net transfer capacities (2 variables): We consider the import net transfer capacities (NTCs) with France and Switzerland obtained from the ENTSO-E transparency platform. These two variables are relevant as Italy (nearly) systematically imports cheaper electricity from France and Switzerland, in the limit of the NTCs. Therefore, variations of the latter directly affect market results – specifically the coupling imports from France and the regular imports (subject to the merit-order) from Switzerland. Note that the other neighboring countries were not considered because of smaller NTCs and the absence of a dominant flow direction.
3.
Day-ahead forecast of wind and solar generation (1 variable): Combined country-level day-ahead forecasts of wind and solar generation obtained from a private data provider (LSEG Data & Analytics, 2025), derived from the European Centre for Medium-Range Weather Forecasts (ECMWF) and Global Forecast System (GFS) weather models.
Note that all these predictors are available at least one hour before the day-ahead market closure at 12:00 and can therefore realistically produce market results forecasts informing trading strategies ”in time”.
Appendix C FPCA results
The dynamic number of supply and demand principal components selected across the test period by our combined elbow-threshold method is pictured on Fig. 10. It turns out this number is rather stable, for both sides with and depending on the period.
Fig. 11 shows the dynamic functional principal components across the test period. The stability suggested by the nearly constant number of selected FPCs is confirmed as we can see that each single FPC roughly captures a fixed mode of variation along time. For supply, FPCs subsequent to the fifth start to be noisy with less consistency. Nevertheless, their is a clear shift in most FPCs, justifying the daily recalibration of FPCA (also motivated by the probable mean shift that is not represented here).
The first five and four supply and demand FPCs can be better interpreted on Fig. 12, which displays the effect a perturbation caused by FPCs on the mean curve for a snapshot of the dynamic FPCA at the middle of the test period. In particular for demand, FPC1 explains the quasi-totality of the variability and concerns the magnitude of the total demanded quantity. The contribution to the total variance beyond FPC1 is negligible. Subsequent FPCs were however selected because the elbow method was in this case less conservative than the threshold method. The additional consideration of their temporal consistency highlighted in 11 suggests their significance and goes against the hypothesis that these FPCs just capture noise. It is likely instead that they could capture the small scale variability in a domain of the curve where the intersection with the supply curve is likely to occur, and that is therefore of interest. For supply, FPC1 measures the overall supply level at any price, while FPC2 contrasts situations where there is higher supply at low price (below 100€/MWh) compared to high price (above 200€/MWh) and vice versa. FPC3 and FPC4 regard the distribution of the prices above (resp. below) 180€/MWh, and similarly contrasts situations when these prices are higher or lower. FPC5 is more difficult to interpret and regards more complex features of the price/volume distribution.
Appendix D Computational efficiency
The computational time requirements for point and probabilistic forecasting models (Table 5 and 6) were evaluated on an Apple M2 Pro chip with a 12-core CPU and 16GB of RAM. Note that the ARX model could be made more efficient by parallelizing model fits across hours or vector components. The remaining models benefit from NumPy’s built-in multithreading for large matrix operations and therefore execute in a multicore fashion.
Price-based point forecasting models are definitely faster as they require fitting only 24 individual regression models, compared with the fits needed for curve-based models. Nevertheless, the overall computational cost remains low across all approaches. For probabilistic forecasting, the price-based normal (N), conformal prediction (CP), and isotonic distributional regression (IDR) models are highly efficient, whereas the quantile regression machine (QRM) is substantially more expensive. However, QRM is the only price-based method that was found competitive against curve-based models, which exhibit comparable computational efficiency.
| Wall time | |
| ARX | 0.2s |
| fARX | 0.6s |
| LEAR | 3s |
| ZST/FPCA-ARX | 2s |
| ZST/FPCA-fARX | 9s |
| ZST/FPCA-VARX | 6s |
| ZST/FPCA-fVARX | 16s |
| Wall time | |
| N | ¡0.1s |
| QRM | 5s |
| CP | ¡0.1s |
| IDR | ¡0.1s |
| ZST/FPCA | 3s |
Appendix E Example of curves and clearing price prediction
Appendix F Reliability and resolution of probabilistic forecasts
Methodology
Probabilistic forecasting performance is composed of a reliability (or calibration) and sharpness (or resolution111111The two concepts are equivalent when probabilistic forecasts have perfect reliability.) component (Nowotarski and Weron, 2018). As any proper scoring rule, the CRPS presents the advantage of simultaneously evaluating both but it can be interesting to understand how models perform on each of these two components. We assess therefore reliability with the probability integral transform (Gneiting and Katzfuss, 2014)
If is actually drawn from , then should be uniformly distributed, which can be verified visually with an histogram (or quantitatively with standard goodness-of-fit statistical tests). Sharpness, when forecasts are reliable, intuitively relates to the capacity of modeling the variability of the response’s uncertainty. A natural way to assess this property is to look at the distribution of a dispersion parameter of the predictive distributions, for instance prediction interval (PI) width or standard deviation. In our experiments, we analyze the joint distribution of the PI width and the magnitude of point forecasts errors to understand if the predicted uncertainty variations are accurate (i.e. larger predicted uncertainty correlates with larger point forecasts errors).
Results
Looking at the calibration and resolution components (Fig. 14 and 15), we notice the poor reliability of the curve-based ZST and price-based fARX-IDR models. The first is strongly over-dispersed while the tails of the second are too light. fARX-N forecasts are reliable but show a poor resolution. The other models show comparable calibration – with a slight under-dispersion and clear skew towards lower prices – but heterogeneous resolutions: low for fARX-CP, intermediate for fARX-QRM and higher for FPCA models.